OpenStudy (anonymous):

Can anyone familiar with series convergence and divergence help me decipher this problem? it's poorly worded and i'm having trouble figuring out exactly what it is asking.

OpenStudy (anonymous):

Let S be the sum of the numbers sin x over those x such that the series: \[\sum_{n=0}^{infinity}(-1)^nn^3(x^3-x^2-8x+7)^n/(3n^4+1)\] converges conditionally. The value of S is?

OpenStudy (anonymous):

Or just, how do you find which x values cause it to converge conditionally.?

OpenStudy (espex):

Thinking out loud here...You can determine if a series converges by finding the limit, much like you would a function, though I'm not certain how the sin(x) fits into it just yet.

OpenStudy (anonymous):

yeah, the sinx part I think just forget about that. it's not really relevant to what the question is really asking. we need values for x such that the series converges conditionally. and that's kind of where I am stuck. How do you test for conditional convergence?

OpenStudy (espex):

Keep in mind that this is exactly where I am currently in my Calc II so I'm not 100%, but the "conditional convergence" could be what "Paul" calls partial sum convergence: http://tutorial.math.lamar.edu/Classes/CalcII/ConvergenceOfSeries.aspx

OpenStudy (anonymous):

Now, since the main topic of this section is the convergence of a series we should mention a stronger type of convergence. A series is said to converge absolutely if also converges. Absolute convergence is stronger than convergence in the sense that a series that is absolutely convergent will also be convergent, but a series that is convergent may or may not be absolutely convergent. In fact if converges and diverges the series is called conditionally convergent.

OpenStudy (anonymous):

that's what he said in there about it. oh but it didn't copy the symbols.

OpenStudy (amistre64):

the values for x on which it converges is called the interval of convergences i beleive

OpenStudy (espex):

So this is part of the power series?

OpenStudy (anonymous):

Right, and you're supposed to test the endpoints or something?

OpenStudy (amistre64):

\[\lim_{n\to\ inf}\frac{(-1)^nn^3(x^3-x^2-8x+7)^n}{3n^4+1}*\frac{3(n-1)^4+1}{(-1)^{n-1}(n-1)^3(x^3-x^2-8x+7)^{n-1}}\]

OpenStudy (amistre64):

\[\lim_{n\to\ inf}\frac{-n^3(x^3-x^2-8x+7)}{3n^4+1}*\frac{3(n-1)^4+1}{(n-1)^3}\] whatever that simplifies to; we pull out the non-n parts

OpenStudy (anonymous):

okay well if I were to say that i'm pretty sure that limit cancels out to |-(x^3-x^2-8x+7)|, what do you suggest I can try next?

OpenStudy (amistre64):

we need to know if the n parts limit to 1 or 0; if 0 then the intervals i -inf to inf if 1 then the radius is from -1 to 1 if thats what we pulled out

OpenStudy (anonymous):

so.. would I try and solve for x values like this.? -1<(x^3-x^2-8x+7)<1

OpenStudy (anonymous):

or can i just use -1 and 1.. or how does that work

OpenStudy (amistre64):

im tryng to get the wolf to verify my idea without going to timedout mode:)

OpenStudy (espex):

Let S be the sum of the numbers sin x over those x such that the series: So does this mean: S = sin(1)/1 + sin(2)/2 + sin(3)/3 ?

OpenStudy (amistre64):

|-x^3+x^2+8x-7)| <= 1 such that 1 is the radius of convergence -1 <= -x^3+x^2+8x-7) <= 1 is our interval of convergence

OpenStudy (amistre64):

-x^3+x^2+8x-7-1 = 0 -x^3+x^2+8x-7+1 = 0

OpenStudy (anonymous):

and then solving that .. would all the solutions for x allow for conditional convergence? and is it less than or equal to? or just less than. I understood that at p = 1 the ratio test is inconclusive..

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=-x%5E3%2Bx%5E2%2B8x-7%3C%3D1+and+-x%5E3%2Bx%5E2%2B8x-7+%3E%3D+-1 there seem to be 3 interval what satisfy those conditions

OpenStudy (anonymous):

So you are saying those endpoints of the three intervals (6 values for x) are the ones that the question is asking for.. the sin(x)/x sum?

OpenStudy (amistre64):

im saying that that is what i come up with for your summation; whether your summation is correct or not, i dunno

OpenStudy (amistre64):

quite frankly, i got no idea if my thought is even correct :/

OpenStudy (anonymous):

Well, now I am just wondering whether maybe the answers are slightly off, because I did this exact same method already and got that exact same three intervals. And I really don't know how else to do it. but something is not working

OpenStudy (amistre64):

is this your wording of the question? or is this the question itself?

OpenStudy (anonymous):

That's the question itself..

OpenStudy (anonymous):

horribly worded.

OpenStudy (amistre64):

well then, if we could get it graphed we would be able to see how well it fits

OpenStudy (amistre64):

it does appear to hit sinx/x in 3 spots

OpenStudy (anonymous):

really? what graph are you using I have no idea how you're getting that?

OpenStudy (amistre64):

i just used the first few sums of the summation you provided; and mapped it with sinx/x

OpenStudy (anonymous):

hmm. so what values were they?

OpenStudy (amistre64):

just using the first 4 sums; we hit in about -3, 1 and 3

OpenStudy (anonymous):

Yeah, i'm not sure. I think we may be interpreting the sinx portion of the question incorrectly. I think it's just asking us to find the x's needed to make that series conditionally convergent. and then just take the sum of sin(x1)/x1+sin(x2/x2) ..etc. depending on how many x values satisfy conditional convergence.

OpenStudy (anonymous):

sin(x2)/x2**

OpenStudy (amistre64):

well, good luck with it :)

OpenStudy (anonymous):

They usually add weird computational portions to questions for our assignments

OpenStudy (anonymous):

yeah, thanks for your help.